Brockman Corporation's earnings per share were $3.50 last year, and its growth rate during the prior 5 years was 10.6% per year. If that growth rate were maintained, how many years would it take for Brockman's EPS to triple?

Select the correct answer.

a. 10.90
b. 13.20
c. 17.80
d. 8.60
e. 15.50
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

(3*3.5)=3.5*(1.106)^n

3=(1.106)^n

Taking log on both sides;

log 3=n*log(1.106)

n=log 3/log(1.106)

=10.90 years(Approx).